OpenStudy (anonymous):

Find the inverse of \[g(x)=\frac{-5+5x}{-3-3x}\]

OpenStudy (saifoo.khan):

y = (-5+5x)/(-3-3x) Solve for x.

OpenStudy (saifoo.khan):

this is tricky..

OpenStudy (anonymous):

I know... is it \[f'(x)=\frac{5-3x}{5+3x}\] Do I leave it as that?

OpenStudy (saifoo.khan):

now did you got that?

OpenStudy (saifoo.khan):

how*

OpenStudy (saifoo.khan):

y = (-5+5x)/(-3-3x) -3y -3xy = -5 + 5x -3y + 5 = 5x + 3xy -3y + 5 = (5+3y)x

OpenStudy (saifoo.khan):

Finally, x = (-3y+5)/(5+3y) Last step: f(x)^-1 = (-3x+5)/(5+3x) phew..

OpenStudy (anonymous):

it's the same thing right, haha... I did the same thing you did...

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (saifoo.khan):

lol. i was in "challenge accepted" mode.. so i didn't saw what you wrote up. sorry. welcome.

OpenStudy (saifoo.khan):

http://www.flickr.com/photos/daviper07/6874536089/

OpenStudy (anonymous):

oh, kml!!!... you all are so funny, making math fun!!!